OpenStudy (anonymous):

A 5-foot-tall person casts an 8-foot shadow. If a vertical pole that supports a basketball hoop is 12 feet high, how long is its shadow?

OpenStudy (anonymous):

5/8=12/x

OpenStudy (anonymous):

i would set a a poportion to solve this

OpenStudy (anonymous):

@ankur504 is right: x = 12*8/5